OpenStudy (anonymous):

In 2009, a city with a population of 679,000 had 2,716 births. What is the number of births per 10,000 in the city? 400 0.4 40 4

OpenStudy (luckycoins888):

@AccessDenied wanna come help

OpenStudy (luckycoins888):

i dont thinkits o.4

OpenStudy (anonymous):

kakarot!

OpenStudy (danjs):

hi

OpenStudy (anonymous):

hi

OpenStudy (danjs):

solve \[\frac{ 679000 }{ 10000 } = \frac{ 2716 }{ x }\]

OpenStudy (anonymous):

u fool kakarot

OpenStudy (anonymous):

i sadly don't know how to solve that problem

OpenStudy (danjs):

\[679000 x = (10000) * (2716)\] \[x = \frac{ 10000*2716 }{ 679000 }\]

OpenStudy (danjs):

40

OpenStudy (anonymous):

thxs

OpenStudy (danjs):

yep